4+2(4x-3(5-x))=2(x+59)
Answer by rfer(10417)

4+2(4x-3(5-x))=2(x+59)
4+2(4x-15+3x)=2x+118
4+14x-30=2x+118
14x-26=2x+118
12x=1144
x=12